If tuition was $10,000 in 2010-2011, how much was it in (a) 2011-2012? (b) 2012-2013? (c) 2013-2014 (See News, p. 132.) Reference page: 132 IN THE NEWS College Tuition Up Agaiin College gets more expensiveagn. Tuition at public fouryear colleges rose 2.9 percent this year, to an average of $8,893 per year. Out-of state students pay an average of $22203 to attend. Private fouryear colleges also saw tuition increases of 2.9 percent-to an average of $30,094. These tuition hikes come on the heels of a 4.5 percent jump in 2012-2013 and 8.5 percent in 2011-2012. Tuition prices have been outpacing general inflation rates for many years. Souro The College Boand ANALYSIS: Tuition increases reduce the real income of students. How much you suffer from inflation depends on what happens to the prices of the products you purchase.
Answer #1

a) For the year 2011-12, tuition hike is 8.5%.

Thus if tuition was $10000 in 2010-11,

tuition in 2011 - 12 = $10000 + 8.5 10000 100

=10000 + 850 = $10850

b)

For the year 2012-13, tuition hike is 4.5%.

Thus if tuition was $10850 in 2011-12,

tuition in 2012 - 13 = $10850 + 4.5 10850 100

=10850 + 488.25 = $11338.25

c)

For this year (2013-14), tuition hike is 2.9%

Hence if tuition was $11338.25 in 2012-13,

tuition in 2013-14 = $11338.25 + 2.9 11338.25 100

= 11338.25 + 328.81 = $ 11667.06